# On the Fontaine-Mazur conjecture for $p=3$

### v3

@@ 4.1. The residually reducible case
2) With Taylor-Wiles primes, we can define the completed homology with auxiliary levels. See [Pan22, Section 4.4] for precise definitions. .
@@ 4.3. The residually irreducible case
Let $R^{\{\mathbf{1}\}}$ be the universal deformation ring parametrizing all deformations of $\bar{\rho}$ which are unipotent when restricted to $I_{F_{v}},v\in\Sigma\setminus\Sigma_{p}$, and with determinant $\chi$ By the universal property, there is a natural surjection $R^{\{\mathbf{1}\}}\twoheadrightarrow\mathbb{T}_{\mathfrak{m}}$. Similar to Definition 4.1.3, we can also define pro-modular and nice primes in this case.
@@
We say a prime of $R^{\{\mathbf{1}\}}$ is pro-modular if it comes from a prime of $\mathbb{T}_{\mathfrak{m}}$. A pro-modular prime $\mathfrak{q}$ of $R^{\{\mathbf{1}\}}$ is nice if
@@
    $R^{\{\mathbf{1}\}}/\mathfrak{q}$ is one-dimensional with characteristic $p$.
@@
Let $\mathfrak{q}\in\operatorname{Spec}R^{\{\mathbf{1}\}}$ be a nice prime. Then the kernel of $(R^{\{\mathbf{1}\}})_{\mathfrak{q}}\to\mathbb{T}_{\mathfrak{q}}$ is nilpotent. Here, we also view it as a prime of $\mathbb{T}_{\mathfrak{m}}$.
@@
If $[F:\mathbb{Q}]\geq 4|\Sigma\setminus\Sigma_{p}|+2$, then the kernel of $R^{\{\mathbf{1}\}}\to\mathbb{T}_{\mathfrak{m}}$ is nilpotent.
@@ 5.1. Setting
By Grunwald-Wang’s theorem (see [NSW08, Theorem 9.2.8]) and soluble base change, we may suppose the number field $F$ satisfies the following assumptions:
@@ 6.1. The proof
Let $\Sigma$ be a finite set of places containing all the ramified finite places of $\rho$ and let $\Sigma_{p}$ be the set of primes above $p$. By Grunwald-Wang’s theorem ([NSW08, Theorem 9.2.8]) and soluble base change, we may assume the following conditions:
@@ A.1. Banach space representations
Let $\Pi$ be a unitary $K$-Banach space representation of $G$ and $\Theta$ an open bounded $G$-invariant lattice in $\Pi$. We denote by $\Theta^{d}$ its Schikhof dual $\Theta^{d}:=\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(\Theta,\mathcal{O})$ equipped with the topology of pointwise convergence.
@@ A.2. The case $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$
For the first part, see [PT21, Theorem 1.4]. For the second part, [Paš13, Theeorem 1.5] shows the result in the case $p\geq 5$ and the case (2) and (3) for $p\geq 3$. If $p=2$ or $p=3$ and we are in the case (1), it is proved in [Paš16, Theorem 1.3]. If $p=3$ and we are in the case (5), [PT21, Theorem 1.4] shows the isomorphism $Z_{\mathfrak{B}}\cong(R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}})_{\mathrm{tf}}$. Since Proposition 2.3.1 shows that the universal deformation ring $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ is $\mathcal{O}$-torsion free, the result also holds in this case. ∎

### v4

@@ 4.1. The residually reducible case
2) With Taylor-Wiles primes, we can define the completed homology with auxiliary levels. See [Pan22, Section 4.4] for precise definitions.
@@ 4.3. The residually irreducible case
Let $R^{\{\xi_{v}\}}$ be the universal deformation ring parametrizing all deformations of $\bar{\rho}$ satisfying $\operatorname{tr}(\rho)|_{I_{F_{v}}}=\xi_{v}+\xi_{v}^{-1}$ for all $v\in\Sigma\setminus\Sigma_{p}$, and with determinant $\chi$ By the universal property, there is a natural surjection $R^{\{\xi_{v}\}}\twoheadrightarrow\mathbb{T}_{\mathfrak{m}}$. Similar to Definition 4.1.3, we can also define pro-modular and nice primes in this case.
@@
We say a prime of $R^{\{\xi_{v}\}}$ is pro-modular if it comes from a prime of $\mathbb{T}_{\mathfrak{m}}$. A pro-modular prime $\mathfrak{q}$ of $R^{\{\xi_{v}\}}$ is nice if
@@
    $R^{\{\xi_{v}\}}/\mathfrak{q}$ is one-dimensional with characteristic $p$.
@@
Let $\mathfrak{q}\in\operatorname{Spec}R^{\{\xi_{v}\}}$ be a nice prime. Then the kernel of $(R^{\{\xi_{v}\}})_{\mathfrak{q}}\to\mathbb{T}_{\mathfrak{q}}$ is nilpotent. Here, we also view it as a prime of $\mathbb{T}_{\mathfrak{m}}$.
@@
If $[F:\mathbb{Q}]\geq 4|\Sigma\setminus\Sigma_{p}|+2$, then the kernel of $R^{\{\xi_{v}\}}\to\mathbb{T}_{\mathfrak{m}}$ is nilpotent.
@@ 5.1. Setting
By Grunwald-Wang’s theorem (see [AT68, Theorem 5, page 80]) and abelian base change, we may suppose the number field $F$ satisfies the following assumptions:
@@ 5.4. The non-generic reducible case
5.5. Potential pro-modularity
In this subsection, we conclude our potential pro-modularity argument.
Let $F$ be an abelian totally real field in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$ and let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$. Let $\bar{\chi}:G_{F,\Sigma}\to\mathbb{F}^{\times}$ be a continuous odd character and we suppose it can be extended to a character of $G_{\mathbb{Q}}$. Assume $\chi:G_{F,\Sigma}\to\mathcal{O}^{\times}$ is a de Rham character as a lifting of $\bar{\chi}$.
Theorem 5.5.1.
Assume that for any place $v|p$, $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$. Then there exists an abelian totally real extension $F_{1}/F$ of even degree such that $p$ splits completely in $F_{1}$ and for any irreducible component of the universal pseudo-deformation ring $R_{F_{1}}^{\mathrm{ps}}$ (parametrizing all pseudo-deformations of $1+\bar{\chi}|_{G_{F_{1}}}$ unramified out places lying above $\Sigma$ with fixed determinant $\chi$) of dimension at least $1+2[F_{1}:\mathbb{Q}]$, it is pro-modular of dimension $1+2[F_{1}:\mathbb{Q}]$.
Proof.
For the case $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1},\omega^{-1}$ for any $v|p$ and $p\geq 5$, the proof is a combination of [Zha24, Theorem 1.0.2] and Grunwald-Wang’s theorem [AT68, Theorem 5, page 80].
For the case $\bar{\chi}|_{G_{F_{v}}}=\omega^{-1}$ and $p=3$, it is proved in this section. Note that in each case of this section, we conclude that the irreducible component $C$ is pro-modular. For the dimension of $C$, from Step 4 of the proof of Theorem 4.1.4, we can find that if $C$ contains a nice prime, then it is of dimension $1+2[F_{1}:\mathbb{Q}]$, and our result follows from Lemma 5.1.2.
For the case $\bar{\chi}|_{G_{F_{v}}}=\omega^{-1}$ and $p\geq 5$, the proof is almost the same as the case $p=3$. We only need to replace the results in Section 2.3 by [Pan22, Lemma 7.4.4 & Lemma 7.4.8]. ∎
We fix an abelian extension $F_{1}/F$ satisfying the properties in Theorem 5.5.1. Similarly as in Section 4.1, we can define the big Hecke algebra $\mathbb{T}_{\psi}$ (without the character $\xi$), and from our construction of $F_{1}$ and the proof of Theorem 5.5.1, we know that there exists a maximal ideal $\mathfrak{m}$ of $\mathbb{T}_{\psi}$ corresponding to the pseudo-representation $1+\bar{\chi}$. In other words, the “pro-modularity” of a prime of $R_{F_{1}}^{\mathrm{ps}}$ here means that it comes from a prime of $\mathbb{T}_{\mathfrak{m}}:=(\mathbb{T}_{\psi})_{\mathfrak{m}}$.
Proposition 5.5.2.
The big Hecke algebra $\mathbb{T}_{\mathfrak{m}}$ is equidimensional of dimension $1+2[F_{1}:\mathbb{Q}]$.
Proof.
It follows from the previous theorem and Corollary 3.3.10. ∎
@@ 6.1. The proof
Let $\Sigma$ be a finite set of places containing all the ramified finite places of $\rho$ and let $\Sigma_{p}$ be the set of primes above $p$. By [Gee22, Fact 4.27] and soluble base change, we may assume the following conditions:
@@ A.1. Banach space representations
Let $\Pi$ be a unitary $K$-Banach space representation of $G$ and $\Theta$ an open bounded $G$-invariant lattice in $\Pi$. We denote by $\Theta^{d}$ its Schikhof dual $\Theta^{d}:=\operatorname{Hom}_{\mathcal{O}}(\Theta,\mathcal{O})$ equipped with the topology of pointwise convergence.
@@ A.2. The case $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$
For the first part, see [PT21, Theorem 1.4]. For the second part, [Paš13, Theorem 1.5] shows the result in the case $p\geq 5$ and the case (2) and (3) for $p\geq 3$. If $p=2$ or $p=3$ and we are in the case (1), it is proved in [Paš16, Theorem 1.3]. If $p=3$ and we are in the case (5), [PT21, Theorem 1.4] shows the isomorphism $Z_{\mathfrak{B}}\cong(R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}})_{\mathrm{tf}}$. Since Proposition 2.3.1 shows that the universal deformation ring $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ is $\mathcal{O}$-torsion free, the result also holds in this case. ∎